### v1

@@ Abstract
Keywords:
Hidden Markov Model Probabilistic Graphical Models Language Models.
@@ 3 Inference in Top-$p$ Hidden Markov Models
$v\in[0,1]^{d},p\in(0,1]$triangleright Input probability distribution and $p$ value
@@
$idx\leftarrow\text{sort\_indices\_by\_value\_desc}(v)$triangleright create sorted order of $v$
@@
for $j$ in $idx$ do $\triangleright$ $j$ is an index for $v$ and $v^{\prime}$
@@
     $v^{\prime}[j]\leftarrow v[j]$triangleright take events …
@@
$v^{\prime}\leftarrow\frac{v^{\prime}}{sum}$triangleright scale to sum to one
@@ 4.1 Approximation Error in a Single Time Step
We show that the error introduced by a top-$p$ distribution is at most $1-p$:
@@ 4.2 Overall Approximation Error
Boyen and Koller show for the Kullback-Leibler divergence that the error in each time step is at most $\frac{\epsilon}{\gamma}$, where $\gamma$ is a parameter inherent to the specific HMM and $\epsilon$ is the error incurred in each time step [5]. The parameter $\gamma$ of an HMM tells the minimum probability mass on which two different state distributions agree on after a single time step. While the bounded error result is promising, we cannot use their theorems directly, because of their use of the Kullback-Leibler divergence, which is not applicable in our case. The Kullback-Leibler divergence divides by probabilities and we set probabilities to zero, rendering the Kullback-Leibler divergence useless in our case. Therefore, we use the key idea from Boyen and Koller [5] to bound the error our top-$p$ HMM in terms of our used total variation distance. We start by defining the parameter $\gamma$ and continue by proving that an error incurred in one time step diminishes by a factor of $1-\gamma$ over the next time steps. Then, given that our error increases at most by $1-p$ in each time step (c.f. Corollary 1), we build a geometric series and prove that our overall approximation error in terms of total variation distance is at most $\frac{1-p}{\gamma}$ in each time step.
@@
We fix $\varphi$ and $\psi$. Using Lemma 1, let $\tilde{\varphi}$ and $\tilde{\psi}$ denote the respective distributions in the intermediate step. We show that our claim holds by showing the error $\delta(\tilde{\varphi},\tilde{\psi})=(1-\gamma)\delta(\varphi,\psi)$ in the intermediate step and the error $\delta(\varphi^{\prime},\psi^{\prime})\leq\delta(\tilde{\varphi},\tilde{\psi})$ in the next time step. The last inequality holds because the total variation is invariant to linear transformations.
@@
$\delta(\tilde{\varphi},\tilde{\psi})$ $=\frac{1}{2}\sum_{i}\left\lvert\tilde{\varphi}(\omega_{i})-\tilde{\psi}(\omega_{i})\right\rvert+\frac{1}{2}\left\lvert\tilde{\varphi}(c)-\tilde{\psi}(c)\right\rvert$ (17)
@@
At this point, we have already shown that we (i) introduce an approximation error of $1-p$ when using the top-$p$ distribution once (c.f. Theorem 4.1), (ii) the error introduced in one time step diminishes by a factor of $1-\gamma$ for each future time step (c.f. Theorem 4.2), which is $(1-p)(1-\gamma)^{k}$ for time step $k$, and (iii) in each time step we only add an additional error of at most $1-p$ (c.f. Corollary 1). We combine these results in our overall approximation guarantee by analyzing the induced geometric series:
@@
In time step $k$, because of the triangle inequality, the error consists of the error from time step $k-1$ and an additional error of at most $1-p$. However, the error of time step $k-1$ is, by Theorem 4.2, diminished by the factor $1-\gamma$. Therefore, the error in time step $k$ is
@@ 5 Evaluation
Sparsity We define the sparsity as the ratio of zero-entries in the transition matrix. In the Bell HMM, all three top-$p$ approaches lead to a sparsity of above $0.99$. In the Uniform HMM, the sparsity is approximately $1-p$. And in the LM HMM, the sparsity is for all three top-$p$ greater than $0.97$.

### v2

@@ 3 Inference in Top-$p$ Hidden Markov Models
$v\in[0,1]^{d},p\in(0,1]$ $\triangleright$ Input probability distribution and $p$ value
@@
$idx\leftarrow\text{sort\_indices\_by\_value\_desc}(v)$ $\triangleright$ create sorted order of $v$
@@
for $j$ in idxidx do $\triangleright$ $j$ is an index for $v$ and $v^{\prime}$
@@
  $v^{\prime}[j]\leftarrow v[j]$ $\triangleright$ take events …
@@
$v^{\prime}\leftarrow\frac{v^{\prime}}{sum}$ $\triangleright$ scale to sum to one
@@ 4.1 Approximation Error in a Single Time Step
We show that the error introduced by a top-$p$ distribution is at most 1p1-p:
@@ 4.2 Overall Approximation Error
Boyen and Koller show for the Kullback-Leibler divergence that the error in each time step is at most $\frac{\epsilon}{\gamma}$, where $\gamma$ is a parameter inherent to the specific HMM and $\epsilon$ is the error incurred in each time step [5]. The parameter $\gamma$ of an HMM tells the minimum probability mass on which two different state distributions agree on after a single time step. While the bounded error result is promising, we cannot use their theorems directly, because of their use of the Kullback-Leibler divergence, which is not applicable in our case. The Kullback-Leibler divergence divides by probabilities and we set probabilities to zero, rendering the Kullback-Leibler divergence useless in our case. Therefore, we use the key idea from Boyen and Koller [5] to bound the error our top-$p$ HMM in terms of our used total variation distance. We start by defining the parameter $\gamma$ and continue by proving that an error incurred in one time step diminishes by a factor of $1-\gamma$ over the next time steps. Then, given that our error increases at most by 1p1-p in each time step (c.f. Corollary 1), we build a geometric series and prove that our overall approximation error in terms of total variation distance is at most $\frac{1-p}{\gamma}$ in each time step.
@@
[…]
@@
[…]
@@
At this point, we have already shown that we (i) introduce an approximation error of 1p1-p when using the top-$p$ distribution once (c.f. Theorem 4.1), (ii) the error introduced in one time step diminishes by a factor of $1-\gamma$ for each future time step (c.f. Theorem 4.2), which is $(1-p)(1-\gamma)^{k}$ for time step $k$, and (iii) in each time step we only add an additional error of at most 1p1-p (c.f. Corollary 1). We combine these results in our overall approximation guarantee by analyzing the induced geometric series:
@@
In time step $k$, because of the triangle inequality, the error consists of the error from time step k1k-1 and an additional error of at most 1p1-p. However, the error of time step k1k-1 is, by Theorem 4.2, diminished by the factor $1-\gamma$. Therefore, the error in time step $k$ is
@@ 5 Evaluation
Sparsity We define the sparsity as the ratio of zero-entries in the transition matrix. In the Bell HMM, all three top-$p$ approaches lead to a sparsity of above $0.99$. In the Uniform HMM, the sparsity is approximately 1p1-p. And in the LM HMM, the sparsity is for all three top-$p$ greater than $0.97$.